An individual has a utility function given by U = x1x2 Marginal Rate of Substitution is –x2/(x1) and therefore the Demand function for good 1 is x1= m/(2P1) Assume m=$42, P1=$1, P2=$1 (m=income, P1 is the price of good 1 , P2 is the price of good 2) Calculate the quantity of good one in the optimal choice bundle (x1A)
Answer #1

The demand function of good 1 is x1 = m/2p1 = 42/2(1) = 21

good one in the optimal choice bundle (x1A)= 21